OpenStudy (anonymous):

Captain Kirk is on an important space mission to bring a Klingon ambassador to space outpost 13. He must travel 3000. km in less than 30.0 min, or else war will be declared between the Klingons and humans. After traveling exactly 27.0 min at a steady pace, he sees that there are still 381.0 km left to go. His ship must then accelerate at 7.167 m/s2 for how many seconds in order to achieve the desired time?

OpenStudy (anonymous):

speed with which he has travelled till 27 minutes is equal to v=1782/27*60= 1100 m/s distance to be covered = 218000m s=ut + 1/2 at*t 218000=198000+16200a 20000=16200a a=20000/16200 =1.2345678